# Exploring Big Five Personality and AI Capability Effects in LLM-Simulated Negotiation Dialogues

### v2

@@ 3.2 Evaluation Measures
Secret Keeping Extent to which agent maintains secrecy of private information or intentions [--10, 0].
Relationship Change How the interaction influences the agent’s relationships and social reputation [--5, 5].
Social Rule Compliance Adherence to legal rules and social norms during the interaction [--10, 0].
Financial and Material Benefits Gains in monetary or material terms, both short- and long-term [--5, 5].
@@ Appendix C Trait Variation Prompt
{minted}
python credibility_persona = ”High_Transparency-High_Competence-High_Adaptability”: ”Task_Assignment”: ”Delegates tasks with clear explanations, leveraging high competence and adaptability to adjust to evolving needs and challenges.”, ”Interaction”: ”Engages openly with team members, sharing knowledge and adapting interactions based on feedback and changing circumstances.”, ”Communication”: ”Communicates transparently and expertly, adapting messages to ensure clarity and relevance for various situations and audiences.”, ”Planning”: ”Involves the team in detailed, transparent planning processes, with strategies that adapt to new information and changing conditions.”, ”Leadership”: ”Leads with high transparency and adaptability, using expertise to navigate changes and inspire confidence and flexibility within the team.”, ”Individual_Role”: ”Known for a high level of openness, skill, and flexibility, significantly contributing to team success by adapting to dynamic environments.” , ”High_Transparency-High_Competence-Low_Adaptability”: ”Task_Assignment”: ”Assigns tasks with clear and competent guidance but may struggle to adjust plans or strategies in response to unforeseen changes.”, ”Interaction”: ”Maintains open communication and provides expert input, though may not easily adapt interactions to rapidly changing team dynamics or feedback.”, ”Communication”: ”Communicates effectively and transparently, but may find it challenging to modify communication styles or approaches as situations evolve.”, ”Planning”: ”Creates detailed plans with clear transparency and high competence, but may have difficulty adapting strategies if new information or changes arise.”, ”Leadership”: ”Leads with clarity and expertise, though adaptability might be limited, potentially affecting the ability to respond effectively to unexpected changes.”, ”Individual_Role”: ”Provides high-quality and transparent input but may need to improve flexibility to better handle evolving situations.” , ”High_Transparency-Low_Competence-High_Adaptability”: ”Task_Assignment”: ”Delegates tasks with openness and clarity but may lack the expertise needed for effective execution, while adapting to team needs and feedback.”, ”Interaction”: ”Engages openly with team members, adapting interactions based on feedback, though might not offer deep or technically sound guidance due to lower competence.”, ”Communication”: ”Communicates transparently and adjusts messaging based on context and feedback, though may lack depth and technical detail in explanations.”, ”Planning”: ”Shares planning processes openly and adapts strategies based on new information, though plans may lack the necessary competence for optimal execution.”, ”Leadership”: ”Promotes transparency and flexibility but may struggle with providing expert guidance, requiring continuous adaptation to improve effectiveness.”, ”Individual_Role”: ”Creates an open and adaptable environment but needs to bolster competence to enhance overall effectiveness and contribution.” , ”High_Transparency-Low_Competence-Low_Adaptability”: ”Task_Assignment”: ”Assigns tasks with clear instructions but struggles with effective execution due to low competence and adaptability, providing minimal updates.”, ”Interaction”: ”Interacts transparently but may be rigid and less responsive to feedback or changing conditions, impacting support and team dynamics.”, ”Communication”: ”Communicates clearly but may lack depth and flexibility, leading to incomplete or inadequate guidance due to limited expertise and adaptability.”, ”Planning”: ”Shares planning details openly but with limited effectiveness and adaptability, resulting in suboptimal strategies and execution challenges.”, ”Leadership”: ”Demonstrates transparency but struggles with both competence and adaptability, leading to less effective leadership and team direction.”, ”Individual_Role”: ”Known for clear but ineffective communication and lack of adaptability, requiring significant improvement in skill and flexibility for effective contribution.” , ”Low_Transparency-High_Competence-High_Adaptability”: ”Task_Assignment”: ”Delegates tasks effectively based on high competence and adaptability but with limited transparency in updates or rationale.”, ”Interaction”: ”Engages positively with team members while adapting interactions based on changing needs, though may not share all relevant information.”, ”Communication”: ”Provides knowledgeable input and adjusts communication style as needed, though might not be fully transparent about processes or details.”, ”Planning”: ”Develops effective and adaptable plans but keeps details and rationale guarded, potentially impacting overall team alignment and understanding.”, ”Leadership”: ”Leads with strong skill and adaptability but maintains some level of secrecy, affecting team trust and cohesion despite effective execution.”, ”Individual_Role”: ”Demonstrates high competence and flexibility but may need to increase transparency to enhance overall team effectiveness and collaboration.” , ”Low_Transparency-High_Competence-Low_Adaptability”: ”Task_Assignment”: ”Assigns tasks with high competence but limited transparency and adaptability, resulting in unclear guidance and difficulty responding to changes.”, ”Interaction”: ”Interacts with caution and minimal openness, providing skilled support but struggling to adapt interactions based on team feedback or changes.”, ”Communication”: ”Communicates authoritatively but with limited transparency, and may struggle to adjust messages based on evolving needs or contexts.”, ”Planning”: ”Creates detailed plans with high expertise but lacks adaptability and transparency, leading to potential gaps in team understanding and responsiveness.”, ”Leadership”: ”Leads with high skill but limited adaptability and openness, which may impact team cohesion and effectiveness despite competent execution.”, ”Individual_Role”: ”Known for high competence but requires improvement in transparency and adaptability to fully support team dynamics and responsiveness.” , ”Low_Transparency-Low_Competence-High_Adaptability”: ”Task_Assignment”: ”Delegates tasks with minimal competence and transparency but shows high adaptability in adjusting approaches based on team feedback and changes.”, ”Interaction”: ”Engages with team members in a flexible manner but may lack depth in technical guidance and provide limited information.”, ”Communication”: ”Communicates with adaptability but limited clarity and expertise, leading to potential misunderstandings and ineffective guidance.”, ”Planning”: ”Plans with high adaptability but minimal transparency and competence, resulting in unclear and potentially ineffective strategies.”, ”Leadership”: ”Demonstrates flexibility and responsiveness but struggles with both transparency and skill, affecting overall leadership effectiveness.”, ”Individual_Role”: ”Creates an adaptable environment but requires significant improvement in competence and transparency to enhance overall effectiveness.” , ”Low_Transparency-Low_Competence-Low_Adaptability”: ”Task_Assignment”: ”Assigns tasks with reluctance and minimal effectiveness, lacking competence, transparency, and adaptability, resulting in poor outcomes.”, ”Interaction”: ”Interacts in a guarded manner with limited information sharing and adaptability, providing minimal support and demonstrating low skill.”, ”Communication”: ”Shares minimal and unclear information, leading to confusion and ineffective communication within the team due to low competence and flexibility.”, ”Planning”: ”Plans with minimal effectiveness and adaptability, resulting in unclear strategies and challenges in execution due to low competence and transparency.”, ”Leadership”: ”Struggles with leadership due to low trust, transparency, competence, and adaptability, leading to poor team dynamics and performance.”, ”Individual_Role”: ”Considered ineffective and uncommunicative, requiring substantial improvement across transparency, competence, and adaptability.”

### v3

@@ 3.2 Evaluation Measures
Secret Keeping Extent to which agent maintains secrecy of private information or intentions [-10, 0].
Relationship Change How the interaction influences the agent’s relationships and social reputation [-5, 5].
Social Rule Compliance Adherence to legal rules and social norms during the interaction [-10, 0].
Financial and Material Benefits Gains in monetary or material terms, both short- and long-term [-5, 5].
@@ Appendix C Trait Variation Prompt
credibility_persona = {
"High_Transparency-High_Competence-High_Adaptability": {
"Task_Assignment": "Delegates tasks with clear explanations, leveraging high competence and adaptability to adjust to evolving needs and challenges.",
"Interaction": "Engages openly with team members, sharing knowledge and adapting interactions based on feedback and changing circumstances.",
"Communication": "Communicates transparently and expertly, adapting messages to ensure clarity and relevance for various situations and audiences.",
"Planning": "Involves the team in detailed, transparent planning processes, with strategies that adapt to new information and changing conditions.",
"Leadership": "Leads with high transparency and adaptability, using expertise to navigate changes and inspire confidence and flexibility within the team.",
"Individual_Role": "Known for a high level of openness, skill, and flexibility, significantly contributing to team success by adapting to dynamic environments."
"High_Transparency-High_Competence-Low_Adaptability": {
"Task_Assignment": "Assigns tasks with clear and competent guidance but may struggle to adjust plans or strategies in response to unforeseen changes.",
"Interaction": "Maintains open communication and provides expert input, though may not easily adapt interactions to rapidly changing team dynamics or feedback.",
"Communication": "Communicates effectively and transparently, but may find it challenging to modify communication styles or approaches as situations evolve.",
"Planning": "Creates detailed plans with clear transparency and high competence, but may have difficulty adapting strategies if new information or changes arise.",
"Leadership": "Leads with clarity and expertise, though adaptability might be limited, potentially affecting the ability to respond effectively to unexpected changes.",
"Individual_Role": "Provides high-quality and transparent input but may need to improve flexibility to better handle evolving situations."
"High_Transparency-Low_Competence-High_Adaptability": {
"Task_Assignment": "Delegates tasks with openness and clarity but may lack the expertise needed for effective execution, while adapting to team needs and feedback.",
"Interaction": "Engages openly with team members, adapting interactions based on feedback, though might not offer deep or technically sound guidance due to lower competence.",
"Communication": "Communicates transparently and adjusts messaging based on context and feedback, though may lack depth and technical detail in explanations.",
"Planning": "Shares planning processes openly and adapts strategies based on new information, though plans may lack the necessary competence for optimal execution.",
"Leadership": "Promotes transparency and flexibility but may struggle with providing expert guidance, requiring continuous adaptation to improve effectiveness.",
"Individual_Role": "Creates an open and adaptable environment but needs to bolster competence to enhance overall effectiveness and contribution."
"High_Transparency-Low_Competence-Low_Adaptability": {
"Task_Assignment": "Assigns tasks with clear instructions but struggles with effective execution due to low competence and adaptability, providing minimal updates.",
"Interaction": "Interacts transparently but may be rigid and less responsive to feedback or changing conditions, impacting support and team dynamics.",
"Communication": "Communicates clearly but may lack depth and flexibility, leading to incomplete or inadequate guidance due to limited expertise and adaptability.",
"Planning": "Shares planning details openly but with limited effectiveness and adaptability, resulting in suboptimal strategies and execution challenges.",
"Leadership": "Demonstrates transparency but struggles with both competence and adaptability, leading to less effective leadership and team direction.",
"Individual_Role": "Known for clear but ineffective communication and lack of adaptability, requiring significant improvement in skill and flexibility for effective contribution."
"Low_Transparency-High_Competence-High_Adaptability": {
"Task_Assignment": "Delegates tasks effectively based on high competence and adaptability but with limited transparency in updates or rationale.",
"Interaction": "Engages positively with team members while adapting interactions based on changing needs, though may not share all relevant information.",
"Communication": "Provides knowledgeable input and adjusts communication style as needed, though might not be fully transparent about processes or details.",
"Planning": "Develops effective and adaptable plans but keeps details and rationale guarded, potentially impacting overall team alignment and understanding.",
"Leadership": "Leads with strong skill and adaptability but maintains some level of secrecy, affecting team trust and cohesion despite effective execution.",
"Individual_Role": "Demonstrates high competence and flexibility but may need to increase transparency to enhance overall team effectiveness and collaboration."
"Low_Transparency-High_Competence-Low_Adaptability": {
"Task_Assignment": "Assigns tasks with high competence but limited transparency and adaptability, resulting in unclear guidance and difficulty responding to changes.",
"Interaction": "Interacts with caution and minimal openness, providing skilled support but struggling to adapt interactions based on team feedback or changes.",
"Communication": "Communicates authoritatively but with limited transparency, and may struggle to adjust messages based on evolving needs or contexts.",
"Planning": "Creates detailed plans with high expertise but lacks adaptability and transparency, leading to potential gaps in team understanding and responsiveness.",
"Leadership": "Leads with high skill but limited adaptability and openness, which may impact team cohesion and effectiveness despite competent execution.",
"Individual_Role": "Known for high competence but requires improvement in transparency and adaptability to fully support team dynamics and responsiveness."
"Low_Transparency-Low_Competence-High_Adaptability": {
"Task_Assignment": "Delegates tasks with minimal competence and transparency but shows high adaptability in adjusting approaches based on team feedback and changes.",
"Interaction": "Engages with team members in a flexible manner but may lack depth in technical guidance and provide limited information.",
"Communication": "Communicates with adaptability but limited clarity and expertise, leading to potential misunderstandings and ineffective guidance.",
"Planning": "Plans with high adaptability but minimal transparency and competence, resulting in unclear and potentially ineffective strategies.",
"Leadership": "Demonstrates flexibility and responsiveness but struggles with both transparency and skill, affecting overall leadership effectiveness.",
"Individual_Role": "Creates an adaptable environment but requires significant improvement in competence and transparency to enhance overall effectiveness."
"Low_Transparency-Low_Competence-Low_Adaptability": {
"Task_Assignment": "Assigns tasks with reluctance and minimal effectiveness, lacking competence, transparency, and adaptability, resulting in poor outcomes.",
"Interaction": "Interacts in a guarded manner with limited information sharing and adaptability, providing minimal support and demonstrating low skill.",
"Communication": "Shares minimal and unclear information, leading to confusion and ineffective communication within the team due to low competence and flexibility.",
"Planning": "Plans with minimal effectiveness and adaptability, resulting in unclear strategies and challenges in execution due to low competence and transparency.",
"Leadership": "Struggles with leadership due to low trust, transparency, competence, and adaptability, leading to poor team dynamics and performance.",
"Individual_Role": "Considered ineffective and uncommunicative, requiring substantial improvement across transparency, competence, and adaptability."